OpenStudy (anonymous):

The perimeter of a octagon is 124 inches. Find the length of each side

OpenStudy (anonymous):

We start with the definition of a regular octagon. Any regular polygon is equilateral. That is, it's sides are of equal length. Also by definition an octagon is an eight sided polygon. Thus a regular octagon has 8 sides of equal lengths. The perimeter of a polygon is the sum of the lengths of each side. This leads to the following equation: 8*(length of a side) = 124 Now we divide both sides by 8 to get length of a side = 124/8 = 15.5 inches

OpenStudy (austinl):

\[8x = 124\] \[x = \frac{124}{8}\]

OpenStudy (austinl):

This is of course assuming that it is a Regular Octagon.